OpenStudy (anonymous):

Solve the equation. 4p+25=6(p-3) - 3(4-3p)

OpenStudy (scorcher219396):

First distribute to get rid of your parentheses 4p+25=6p-18-12+9p Then simplify what you can (combine everything that has the same variable on each side, and everything that has no variables on each side) 4p+25=15p-30 Move your variables to one side, and your constants to the other 4p+25=15p-30 -15p -15p -11p+25=-30 -25 -25 -11p=-55 Then divide by the coefficient in front of your variable (the number in front of your variable) -11p=-55 /-11 /-11 p=-55/-11, then simplify to 5